# Multi-object Data Integration in the Study of Primary Progressive Aphasia

### v1

@@ 1 Introduction
This article is motivated by a clinical application on patients suffering from brain loss due to Primary Progressive Aphasia (PPA), a neurodegenerative disorder (ND) characterized by loss of language ability that shares pathological signatures with Alzheimer’s disease and frontotemporal dementia. Despite the devastating toll PPA effects on a patient’s ability to communicate, it is far less widely studied than other NDs, and as a result, research cohorts contain modest sample sizes. We consider the nonfluent/agrammatic variant of PPA (nfvPPA) characterized by motor speech and grammar loss and left inferior frontal atrophy (Gorno-Tempini et al., 2008). Multi-modal imaging data are available for each of these PPA patients which include: (a) brain network information and (b) brain structural information. Brain networks consider each region of interest (ROI) as a network node and quantifies the connectivity between the pairs of network nodes using functional magnetic resonance imaging (fMRI). Brain structural information is obtained using structural magnetic resonance imaging (sMRI), e.g., gray matter (GM) images over brain volumetric pixels (voxels). Both images are collected on a common brain atlas which segments a human brain into different ROIs, with each ROI containing a number of voxels. A speech rate score is available for all these patients to measure the degree of their motor speech loss due to PPA.
The major scientific goals of the study are twofold, (a) to estimate regression relationships between the speech rate and the two imaging modalities; and (b) to identify ROIs significantly related to the speech rate, and thus infer the ND patterns for PPA.
@@
To this end, the identification of statistically influential ROIs via a principled modeling framework is necessary to yield results that are scientifically interpretable in the lingua franca of well-characterized brain anatomy and function. While many ROI-based inferential techniques are constrained by considering the ROI as the most basic unit of analysis, the PPA data application allows for additional granular inference at the level of the voxel which will provide neuroscientists with the dual ability to frame results at the macro level of the ROI and probe the micro voxel-level signals that drive these results. To our knowledge, this article proposes the first statistical framework that allows for flexibility of macro and micro inference on multi-object imaging data which is urgently needed in the neuroscience community (Calhoun and Sui, 2016). Uncertainty quantification regarding the inference on influential regions becomes crucial since the analysis only involves a limited number of rarely studied PPA patients.
@@
This article addresses the inferential objective by formulating a multi-modal regression framework with the brain network (defined over brain ROIs) and structural images (defined over brain voxels) as two sets of responses and the speech rate as the predictor. A hierarchical Bayesian approach is adopted wherein an integrated prior structure on coefficients of the speech rate corresponding to the structural and network objects (hereon referred to as the structural and network coefficients, respectively) is developed to allow the information in separate image objects to complement and reinforce each other in their relation to the scalar predictor. To elaborate, we begin with a common brain atlas for both image objects to ensure an organizing principle that links together structural and network information via a shared set of ROIs and a group of voxels within each ROI. The integrated prior on network and structural coefficients are constructed respecting the hierarchical constraint which ensures all voxels within an ROI uninfluential if the ROI is un-influential (see Section 3.2). While we do not explicitly make use of the structural information in the GM images by careful spatial modeling of the GM image coefficient, it partially exploits the structural information by imposing the hierarchical constraint in the prior construction. The problem of identifying influential ROIs is cast under a nonlinear variable selection framework, wherein latent activation indicators corresponding to all ROIs are shared among both sets of coefficients to enforce that all voxels from a particular ROI and all network edges connected to that particular ROI have no relationship with the predictor when the activation indicator corresponding to the ROI is zero. As a byproduct of our construction, symmetry and transitivity property (Hoff, 2005) of the un-directed network object is preserved, as discussed in Section 3.2. The prior construction achieves efficient computation, identifies influential ROIs and voxels within an influential ROI which are key to studying neuronal degradation, and produces well-calibrated interval estimates for the multi-modal regression coefficients. Moreover, our framework attaches uncertainty in identifying these ROIs and offers improved inference over regression methods with a single imaging modality.
@@
There is a dearth of principled Bayesian literature addressing the inferential objectives of the motivating application, and our proposal is arguably the first Bayesian multi-object regression approach to answer the inferential questions stated before. We now provide a brief overview of the available literature to contrast them with our proposal. In the course to determine the association between network or structural objects and the speech rate score, the most popular approach estimates the association between each network edge or GM voxel and the language score independently, providing a p-value “map” (Lazar, 2008, 2016), after adjusting for multiple comparisons to identify “significant” ROIs. Such approaches have key disadvantages relative to methods that take into account the integrated impact of all network edges and voxels of the GM images simultaneously, as demonstrated in Section 5.
@@
One can possibly proceed to vectorize both objects and regress them jointly on the speech rate, leading to a high dimensional vector response regression problem. This approach can take advantage of the recent developments in high dimensional multivariate reduced rank sparse regression literature (Rothman et al., 2010; Chen and Huang, 2012; Goh et al., 2017). However, vectorization of both sets of object responses during analysis ignores their individual characteristics (e.g., the symmetry and transitivity in the network) and their interconnections through the hierarchical constraint, and apparently does not allow identification of influential ROIs with uncertainty.
@@
There have been some recent efforts to build regression models with a single-object response (e.g., tensors, networks, or functions), mainly for neuro-imaging data (Guhaniyogi and Spencer, 2021; Spencer et al., 2020; Zeng et al., 2022; Guha and Guhaniyogi, 2021; Morris et al., 2022). While these approaches establish the importance of exploiting the structure and/or network characteristics in image objects for better inference, principled linkages among image objects are not made (e.g., through the hierarchical constraint) and thus inference is made without regard to valuable information shared across these objects. Failure to consider the structure and cross information from multiple images (i.e. unimodal analysis) have generally a negative impact on ND research in terms of bias in estimated effects, statistical inefficiency (Dai and Li, 2021), and sensitivity of results to noise (Calhoun and Sui, 2016).
@@
Our proposed approach is considerably different from the existing statistical literature on multi-modal data integration. In particular, there has been a class of unsupervised multi-modal analysis with methods exploiting structural connectivity information from diffusion tensor imaging (DTI) in the prior construction for the functional connectivity analysis from functional MRI (fMRI) data (Xue et al., 2015). Building on work in sparse canonical correlation analysis (Witten and Tibshirani, 2009; Chen et al., 2013), there is a growing literature in imaging genetics that seeks to model the cross-covariance between a set of high-dimensional genetic features and imaging phenotypes from different modalities (Du et al., 2016, 2018, 2021; Zhang et al., 2022) which occasionally integrates linkage disequilibrium into estimation but does not explicitly model the topology of the imaging data. With regard to supervised learning approaches, nearly all approaches treat multi-modal imaging data as a predictor rather than a response. Xue et al. (2018) proposes regression of disease status on low-frequency fluctuation (fALFF) from resting-state fMRI scans, voxel-based morphometry (VBM) from T1-weighted MRI scans, and fractional anisotropy (FA) from DTI scans. In the same vein Li and Li (2021); Dai and Li (2021) develop frameworks to account for the non-linear association between a scalar response and multi-modal predictors. Peng et al. (2016) utilizes sparse kernel learning with a novel group penalty to integrate information among genetics and multi-modal images to predict Alzheimer’s status. Ma et al. (2023); Guo et al. (2022) both develop Bayesian regression models that predict a single response image from multi-modal image predictors. Recently, there has been a flurry of attention focused on generative deep learning which seeks to generate multi-modal images based on input features including disease status (Suzuki and Matsuo, 2022; Fernandez et al., 2024) for the purpose of data augmentation and synthetic image generation. While these unsupervised and supervised approaches do form linkages among image modalities, they commonly employ regression frameworks with a scalar response and multi-modal imaging predictors. Such a framework presents two major challenges for our application of interest. First, the limited sample size in our study (sample size of 24, see Section 2) necessitates significantly downsizing each image and assuming strong sparsity in the image predictors to achieve statistically meaningful inferences from the high-dimensional regression, which is less appealing to neuroscientists. Second, image voxels often exhibit collinearity, resulting in less reliable inference when used as predictors in high-dimensional regression. We address this issue by performing integrated modeling of structural and network images as a function of a primary predictor of interest while maintaining image topology. Our framework is able to deliver inference on influential ROIs which is the main scientific objective of the clinical study. Notably, all these aforementioned approaches do not naturally offer identification of influential ROIs with uncertainty.
@@
Up to this point, the foundational work in the cognitive neuroscience of PPA (Mandelli et al., 2016; Gorno-Tempini et al., 2011, 2008) has not directly combined structural information relating to neuronal atrophy with network information on brain connectivity. At best, brain networks that underlie language have been identified via an informal fusion of information from multi-object image data via seed-based methods (Battistella et al., 2020). Specifically, several ROIs are identified that display the highest degree of atrophy in structural images (e.g. GM degradation) of PPA patients. These ROIs are then used to define “seed regions” which are used to construct speech networks that quantify the nodes highly connected to the seed regions. While the speech networks identified via this approach are likely important to study PPA, the ad-hoc way of selecting seed regions means that the influence of atrophy in other ROIs may have been overlooked and inference may have been subject to the limitations of unimodal analysis outlined above. Thus, the application of our proposed approach will provide a new range of opportunities anchored in multi-modal image integration to study cognitive language models and quantify the joint role of structural atrophy and functional connectivity in dementia-related language loss.
@@ 2 Clinical Case Study on Nonfluent Primary Progressive Aphasia (nfvPPA)
We focus on a clinical application derived from multi-modal image studies conducted on patients with PPA, an ND that impacts brain networks that support language, destroying an individual’s ability to communicate with loved ones and leading inevitably to death (Gorno-Tempini et al., 2011). Our interest lies in the nonfluent/agrammatic variant of PPA (nfvPPA) characterized by motor speech and grammar loss and left inferior frontal atrophy (Gorno-Tempini et al., 2008). To investigate the neural underpinnings of disruption to motor speech/fluency in nfvPPA patients, clinical images from multiple modalities were collected as detailed below. Only a handful of institutions worldwide collect brain scans on PPA patients for research, and thus our sample size is constrained due to the prevalence of the disease and resource limitations. However, the severity of the disease and its utility as a lens to understand the neuroanatomy of language production provide sufficient clinical motivation to proceed with analysis while practicing sufficient caution in interpretation due to the modest sample size.
@@
Clinical images and language evaluation: PPA being an infrequently phenotyped ND, imaging data is acquired on $24$ nfvPPA patients during the course of clinical research activity. Data is collected from the following imaging modalities: sMRI derived GM values (Figure 1a) which measures the segmentation probability that a voxel contains neuronal cell bodies (Rajapakse et al., 1997), and task-free resting state fMRI to measure brain activation via neuronal oxygen consumption in subjects at rest. All images are registered to the Montreal Neurological Institute (MNI) template space with voxels parcellated into 245 ROIs using the Brainnetome atlas such that images across modalities and subjects can be directly compared and each voxel is nested in an anatomically defined ROI (Figure 1b) (Fan et al., 2016). The choice of brain atlas is an important decision in the analysis pipeline as it impacts subsequent results and interpretations. Based on our collaborators’ prior work studying language in patients with PPA, the Brainnetome atlas is selected since it provides a granular parcellation needed to model the subtle structural and network dynamics of speech fluency.
For each subject, a ‘brain network’ represented by a symmetric adjacency matrix is obtained from the fMRI image by considering rows and columns of this matrix corresponding to different ROIs and entries corresponding to the z-scores obtained by Fisher Z-transforming the Pearson correlation of the average fMRI data of two ROIs (Figure 1c). We focus our analysis of language loss on speech rate, the number of words spoken per minute, a measure of motor speech via a subject’s articulation rate. This speech rate measure is automatically extracted from recorded speech from the Grandfather passage, a 129-word block of text meant to elicit a comprehensive set of phonemes in English (Ogar et al., 2007), via SALT software (https://www.saltsoftware.com/), a software platform used to automatically extract language features from recorded speech.
@@
Scientific question of interest: Language loss in nfvPPA patients is driven by neurodegeneration in the left inferior frontal region but the dual role of structural damage and brain connectivity in language loss is not well characterized (Mandelli et al., 2016). Findings from the prior clinical studies based on individual image modalities have identified brain networks related to motor speech/fluency (Mandelli et al., 2016). One of our primary goals is to extend the study of motor speech/fluency to consider multi-modal image data spanning the whole brain to better understand how structural and network changes simultaneously impact speech ability. To do so, we apply our proposed multi-modal framework to regress sophisticated multi-modal images, specifically GM maps which capture focal neurodegeneration, and fMRI brain connectivity networks which capture disruptions of brain connectivity, on the speech rate score. The next section describes the novel regression framework to achieve these scientific goals.
@@ 3.1 Model Framework
For the $i$th subject, let $y_{i}\in\mathbb{R}$ denote the speech rate measure, ${\boldsymbol{x}}_{i}=(x_{i,1},...,x_{i,H})^{\prime}$ represent the biological and demographic variables and ${\boldsymbol{A}}_{i}$ represent the weighted brain network object. We assume that the network of all subjects is defined on a common set of nodes, with elements of ${\boldsymbol{A}}_{i}$ encoding the strength of network connections between different nodes for the $i$-th subject. In particular, the network object ${\boldsymbol{A}}_{i}$ is expressed in the form of a $P\times P$ matrix with the $(p,p^{\prime})$-th entry of the matrix $a_{i,(p,p^{\prime})}$ signifying the strength of association between the $p$th and $p^{\prime}$th node, where $p,p^{\prime}=1,...,P$ and $P$ is the number of network nodes. This paper specifically focuses on networks that contain no self-relationship, i.e., $a_{i,(p,p)}\equiv 0$, and are un-directed ($a_{i,(p,p^{\prime})}=a_{i,(p^{\prime},p)}$). Such assumptions hold for the data application pertaining to Section 2, where ${\boldsymbol{A}}_{i}$ represents the brain connectome network matrix obtained from the fMRI scan, with each node representing a specific brain ROI and edges signifying correlations between fMRI signals in two regions. Let ${\boldsymbol{g}}_{i,1}$,…,${\boldsymbol{g}}_{i,P}$ denote the $V_{1},...,V_{P}$ dimensional structural objects in regions $\mathcal{R}_{1}$,…,$\mathcal{R}_{P}$, respectively. In the context of our data application, they represent volumetric elements (voxels) of the GM image from the $P$ ROIs. The nested structure of voxels within ROIs provides biologically plausible organization and is instrumental for variable selection and computation as detailed in the upcoming methodological development.
For $i=1,\ldots,n$, we assume that the relationship between the speech rate measure $y_{i}$ varies with every network edge and every GM voxel after accounting for the edge and voxel-specific effect of the biological and demographic covariates, and propose a set of conditionally independent generalized linear models for every network edge and GM voxel, given by
@@
for $v=1,\ldots,V_{p};\>\>p=1,\ldots,P$, where $H_{1}(\cdot)$ and $H_{2}(\cdot)$ are the link functions, $\theta_{p,p^{\prime}}$ is the $(p,p^{\prime})$th element of the $P\times P$ matrix ${\boldsymbol{\Theta}}$, $\beta_{v,p}$ is the $v$th element of the $V_{p}$ dimensional vector of coefficients ${\boldsymbol{\beta}}_{p}$, $\psi_{p,p^{\prime},h}^{a}$ denotes the effect of $x_{i,h}$ on the edge connecting $p$th and $p^{\prime}$th nodes of the network matrix and $\psi_{p,h}^{g}$ determines the effect of $x_{i,h}$ on the GM image at a voxel within $\mathcal{R}_{p}$. Considering the symmetry and zero diagonal constraints in the network object ${\boldsymbol{A}}_{i}$, we set $\theta_{p,p^{\prime}}=\theta_{p^{\prime},p}$ and $\theta_{p,p}=0$, for all $1\leq p<p^{\prime}\leq P$. Additionally, we assume a bilinear structure for network effects for covariates, given by $\psi_{p,p^{\prime},h}^{a}=\psi_{p,h}^{a}\psi_{p^{\prime},h}^{a}$, to reduce the parameter space while allowing a sufficiently flexible structure to address potential confounding bias in our primary predictor of interest (Von Rosen, 2018). When both sets of responses follow a normal linear model with an identity link function, (1) becomes

### v2

@@ 1 Introduction
This article is motivated by a clinical application on patients suffering from brain loss due to Primary Progressive Aphasia (PPA), a neurodegenerative disorder (ND) characterized by loss of language ability that shares pathological signatures with Alzheimer’s disease and frontotemporal dementia. Despite the devastating toll PPA effects on a patient’s ability to communicate, it is far less widely studied than other NDs, and as a result, research cohorts contain modest sample sizes. We consider the nonfluent/agrammatic variant of PPA (nfvPPA) characterized by motor speech and grammar loss and left inferior frontal atrophy (Gorno-Tempini et al., 2008). Multi-modal imaging data are available for each of these PPA patients which include: (a) brain network information and (b) brain structural information. Brain networks consider each region of interest (ROI) as a network node and quantifies the connectivity between the pairs of network nodes using functional magnetic resonance imaging (fMRI). Brain structural information is obtained using structural magnetic resonance imaging (sMRI), e.g., gray matter (GM) images over brain volumetric pixels (voxels). Both images are collected on a common brain atlas which segments a human brain into different ROIs, with each ROI containing a number of voxels. A speech rate score is available for the patients to measure the degree of their motor speech loss due to PPA.
@@
The major scientific goals of the study are twofold, (a) to estimate regression relationships between the speech rate and the two imaging modalities; and (b) to identify ROIs significantly related to the speech rate, and thus infer the ND patterns for PPA. While many ROI-based inferential techniques are constrained by considering the ROI as the most basic unit of analysis, the PPA data application allows for additional granular inference at the level of the voxel which will provide neuroscientists with the dual ability to frame results at the macro level of the ROI and probe the micro voxel-level signals that drive these results. To our knowledge, this article proposes the first statistical framework that allows for flexibility of macro and micro inference on multi-object imaging data which is urgently needed in neuroscience (Calhoun and Sui, 2016). Uncertainty quantification regarding the inference on influential regions becomes crucial since the analysis only involves a limited number of rarely studied PPA patients.
@@
This article addresses the inferential objective by formulating a multi-modal regression framework with the brain network (defined over brain ROIs) and structural images (defined over brain voxels) as two sets of responses and the speech rate as the predictor. A hierarchical Bayesian approach is adopted wherein an integrated prior structure on coefficients of the speech rate corresponding to the structural and network objects (hereon referred to as the structural and network coefficients, respectively) is developed to allow the information in separate image objects to complement and reinforce each other in their relation to the scalar predictor. To elaborate, we begin with a common brain atlas for both image objects to ensure an organizing principle that links together structural and network information via a shared set of ROIs and a group of voxels within each ROI. The integrated prior on network and structural coefficients are constructed respecting the hierarchical constraint which ensures all voxels within an ROI uninfluential if the ROI is un-influential (see Section 3.2). While we do not explicitly make use of the structural information in the GM images by careful spatial modeling of the GM image coefficient, it partially exploits the structural information by imposing the hierarchical constraint in the prior construction. The problem of identifying influential ROIs is cast under a nonlinear variable selection framework, wherein latent activation indicators corresponding to all ROIs are shared among both sets of coefficients to enforce that all voxels from a particular ROI and all network edges connected to that particular ROI have no relationship with the predictor when the activation indicator corresponding to the ROI is zero. As a byproduct of our construction, the symmetry and transitivity properties (Hoff, 2005) of the un-directed network object are preserved, as discussed in Section 3.2. The prior construction achieves efficient computation, identifies influential ROIs and voxels within an influential ROI which are key to studying neuronal degradation, and produces well-calibrated interval estimates for the multi-modal regression coefficients. Moreover, our framework attaches uncertainty in identifying these ROIs and offers improved inference over regression methods with a single imaging modality.
@@
There is a dearth of principled Bayesian literature addressing the inferential objectives of the motivating application, and our proposal is arguably the first Bayesian multi-object regression approach to answer the inferential questions stated before. We now provide a brief overview of the available literature to contrast them with our proposal. In the course to determine the association between network or structural objects and the speech rate score, the most popular approach estimates the association between each network edge or GM voxel and the language score independently, providing a p-value “map” (Lazar, 2008), after adjusting for multiple comparisons to identify “significant” ROIs. Such approaches have key disadvantages relative to methods that take into account the integrated impact of all network edges and voxels of the GM images simultaneously, as demonstrated in Section 5.
@@
One can possibly proceed to vectorize both objects and regress them jointly on the speech rate with high dimensional multivariate reduced rank sparse regression (Chen and Huang, 2012; Goh et al., 2017). However, vectorization of object responses during analysis ignores their individual characteristics (e.g., the symmetry and transitivity in the network) and their interconnections through the hierarchical constraint, and apparently does not allow identification of influential ROIs with uncertainty.
@@
There have been some recent efforts to build regression models with a single-object response (e.g., tensors, networks, or functions), mainly for neuro-imaging data (Spencer et al., 2020; Zeng et al., 2024; Guha and Guhaniyogi, 2021). While these approaches establish the importance of exploiting the structure and/or network characteristics in image objects for better inference, principled linkages among image objects are not made (e.g., through the hierarchical constraint) and thus inference is made without regard to information shared across these objects. Failure to consider the structure and cross information from images have generally a negative impact on ND research in terms of bias in estimated effects, statistical inefficiency (Dai and Li, 2023), and sensitivity of results to noise (Calhoun and Sui, 2016).
@@
Our proposed approach is considerably different from the existing statistical literature on multi-modal data integration. In particular, there has been a class of unsupervised multi-modal analysis with methods exploiting structural connectivity information from diffusion tensor imaging (DTI) in the prior construction for the functional connectivity analysis from functional MRI (fMRI) data (Xue et al., 2015). Building on work in sparse canonical correlation analysis (Witten and Tibshirani, 2009; Chen et al., 2013), there is a growing literature in imaging genetics that seeks to model the cross-covariance between a set of high-dimensional genetic features and imaging phenotypes from different modalities (Du et al., 2018; Zhang et al., 2022) which occasionally integrates linkage disequilibrium into estimation but does not explicitly model the topology of the imaging data. With regard to supervised learning approaches, nearly all approaches treat multi-modal imaging data as a predictor rather than a response. Xue et al. (2018) proposes regression of disease status on low-frequency fluctuation (fALFF) from resting-state fMRI scans, voxel-based morphometry (VBM) from T1-weighted MRI scans, and fractional anisotropy (FA) from DTI scans. In the same vein Li and Li (2021); Dai and Li (2023) develop frameworks to account for the non-linear association between a scalar response and multi-modal predictors. Peng et al. (2016) utilizes sparse kernel learning with a novel group penalty to integrate information among genetics and multi-modal images to predict Alzheimer’s status. Ma et al. (2023); Guo et al. (2022) both develop Bayesian regression models that predict a single response image from multi-modal image predictors. Recently, there has been a flurry of attention focused on generative deep learning which seeks to generate multi-modal images based on input features including disease status (Suzuki and Matsuo, 2022; Fernandez et al., 2024) for the purpose of data augmentation and synthetic image generation. These approaches commonly employ regression frameworks with a scalar response and multi-modal imaging predictors which presents two major challenges for our application of interest. First, the limited sample size in our study (sample size of 24, see Section 2) necessitates significantly downsizing each image and assuming strong sparsity in the image predictors to achieve statistically meaningful inferences from the high-dimensional regression, which is less appealing to neuroscientists. Second, image voxels often exhibit collinearity, resulting in less reliable inference when used as predictors in high-dimensional regression. We address this issue by performing integrated modeling of structural and network images as a function of a primary predictor of interest. Our framework is able to deliver inference on influential ROIs which is the main scientific objective of the clinical study. Notably, all these aforementioned approaches do not naturally offer identification of influential ROIs with uncertainty.
@@
Up to this point, the foundational work in the cognitive neuroscience of PPA (Mandelli et al., 2016; Gorno-Tempini et al., 2011, 2008) has not directly combined structural information relating to neuronal atrophy with network information on brain connectivity. At best, brain networks that underlie language have been identified via an informal fusion of information from multi-object image data via seed-based methods (Battistella et al., 2020). Specifically, several ROIs are identified that display the highest degree of atrophy in structural images (e.g. GM degradation) of PPA patients. These ROIs are then used to define “seed regions” which are used to construct speech networks that quantify the nodes highly connected to the seed regions. While the speech networks identified via this approach are likely important to study PPA, the ad-hoc way of selecting seed regions means that the influence of atrophy in other ROIs may have been overlooked and inference may have been subject to the limitations of unimodal analysis.
@@ 2 Clinical Case Study on Nonfluent Primary Progressive Aphasia (nfvPPA)
We focus on a clinical application derived from multi-modal image studies conducted on patients with PPA, an ND that impacts brain networks that support language (Gorno-Tempini et al., 2011). Our interest lies in the nonfluent/agrammatic variant of PPA (nfvPPA) characterized by motor speech and grammar loss and left inferior frontal atrophy (Gorno-Tempini et al., 2008). To investigate the neural underpinnings of disruption to motor speech/fluency in nfvPPA patients, clinical images from multiple modalities were collected as detailed below. Only a handful of institutions worldwide collect brain scans on PPA patients for research, and thus our sample size is constrained due to the prevalence of the disease and resource limitations. However, the utility of the disease as a lens to understand the neuroanatomy of language production provide sufficient clinical motivation to proceed with analysis while practicing sufficient caution in interpretation due to the modest sample size.
@@
Clinical images and language evaluation: PPA being an infrequently phenotyped ND, imaging data is acquired on $24$ nfvPPA patients during the course of clinical research activity. Data is collected from the following imaging modalities: sMRI derived GM values (Figure 1a) which measures the segmentation probability that a voxel contains neuronal cell bodies (Rajapakse et al., 1997), and task-free resting state fMRI to measure brain activation via neuronal oxygen consumption in subjects at rest. All images are registered to the Montreal Neurological Institute (MNI) template space with voxels parcellated into 245 ROIs using the Brainnetome atlas such that images across modalities and subjects can be directly compared and each voxel is nested in an anatomically defined ROI (Figure 1b) (Fan et al., 2016). The choice of brain atlas is an important decision in the analysis pipeline as it impacts subsequent results and interpretations. Based on our collaborators’ prior work studying language in patients with PPA, the Brainnetome atlas is selected since it provides a granular parcellation needed to model the subtle structural and network dynamics of speech fluency. For each subject, a ‘brain network’ represented by a symmetric adjacency matrix is obtained from the fMRI image by considering rows and columns of this matrix corresponding to different ROIs and entries corresponding to the z-scores obtained by Fisher Z-transforming the Pearson correlation of the average fMRI data of two ROIs (Figure 1c). We focus our analysis of language loss on speech rate, the number of words spoken per minute, a measure of motor speech via a subject’s articulation rate. This speech rate measure is automatically extracted from recorded speech from the Grandfather passage, a 129-word block of text meant to elicit a comprehensive set of phonemes in English (Ogar et al., 2007), via SALT software (https://www.saltsoftware.com/), a software platform used to automatically extract language features from recorded speech.
@@
Scientific question of interest: Language loss in nfvPPA patients is driven by neurodegeneration in the left inferior frontal region but the dual role of structural damage and brain connectivity in language loss is not well characterized (Mandelli et al., 2016). Findings from the prior clinical studies based on individual image modalities have identified brain networks related to motor speech/fluency (Mandelli et al., 2016). One of our primary goals is to extend the study of motor speech/fluency to consider multi-modal image data spanning the whole brain to better understand how structural and network changes simultaneously associate with speech ability. This scientific question is motivated by findings in PPA and other NDs that observe alignment of disruptions in functional connectivity and structural atrophy (Johnson et al., 2000; Battistella et al., 2019; Brown et al., 2019; Raj et al., 2015) and thus our inference inherently focuses identifying and quantifying these co-localized effects and their relation to language loss. The exact mechanism underlying this co-localization is not known, and may not fully explain language loss in nfvPPA patients, but characterizing and measuring these co-localized effects is only possible by analyzing multiple image modalities and is needed to advance understanding of disease progression and presentation. To do so, we apply our proposed multi-modal framework to regress sophisticated multi-modal images, specifically GM maps which capture focal neurodegeneration, and fMRI brain connectivity networks which capture disruptions of brain connectivity, on the speech rate score. The next section describes the novel regression framework to achieve these scientific goals.
@@ 3.1 Model Framework
For the $i$th subject, let $y_{i}\in\mathbb{R}$ denote the speech rate measure, ${\boldsymbol{x}}_{i}=(x_{i,1},...,x_{i,H})^{\prime}$ represent the biological and demographic variables and ${\boldsymbol{A}}_{i}$ represent the weighted brain network object. We assume that the network of all subjects is defined on a common set of nodes, with elements of ${\boldsymbol{A}}_{i}$ encoding the strength of network connections between different nodes for the $i$-th subject. In particular, the network object ${\boldsymbol{A}}_{i}$ is expressed in the form of a $P\times P$ matrix with the $(p,p^{\prime})$-th entry of the matrix $a_{i,(p,p^{\prime})}$ signifying the strength of association between the $p$th and $p^{\prime}$th node, where $p,p^{\prime}=1,...,P$ and $P$ is the number of network nodes. This paper specifically focuses on networks that contain no self-relationship, i.e., $a_{i,(p,p)}\equiv 0$, and are un-directed ($a_{i,(p,p^{\prime})}=a_{i,(p^{\prime},p)}$). Such assumptions hold for the data application pertaining to Section 2, where ${\boldsymbol{A}}_{i}$ represents the brain connectome network matrix obtained from the fMRI scan, with each node representing a specific brain ROI and edges signifying correlations between fMRI signals in two regions. Let ${\boldsymbol{g}}_{i,1}$,…,${\boldsymbol{g}}_{i,P}$ denote the $V_{1},...,V_{P}$ dimensional structural objects in regions $\mathcal{R}_{1}$,…,$\mathcal{R}_{P}$, respectively. In the context of our data application, they represent voxels of the GM image from the $P$ ROIs.
For $i=1,\ldots,n$, we assume that the relationship between the speech rate $y_{i}$ varies with every network edge and every GM voxel after accounting for the edge and voxel-specific effect of the biological and demographic covariates, and propose conditionally independent generalized linear models for every network edge and GM voxel, given by
@@
for $v=1,\ldots,V_{p};\>\>p=1,\ldots,P$, where $H_{1}(\cdot)$ and $H_{2}(\cdot)$ are the link functions, $\theta_{p,p^{\prime}}$ is the $(p,p^{\prime})$th element of the $P\times P$ matrix ${\boldsymbol{\Theta}}$, $\beta_{v,p}$ is the $v$th element of the $V_{p}$ dimensional vector of coefficients ${\boldsymbol{\beta}}_{p}$, $\psi_{p,p^{\prime},h}^{a}$ denotes the effect of $x_{i,h}$ on the edge connecting $p$th and $p^{\prime}$th nodes of the network matrix and $\psi_{p,h}^{g}$ determines the effect of $x_{i,h}$ on the GM image at a voxel within $\mathcal{R}_{p}$. Considering the symmetry and zero diagonal constraints in the network object ${\boldsymbol{A}}_{i}$, we set $\theta_{p,p^{\prime}}=\theta_{p^{\prime},p}$ and $\theta_{p,p}=0$, for all $1\leq p<p^{\prime}\leq P$. Additionally, we assume a bilinear structure for network effects for covariates, given by $\psi_{p,p^{\prime},h}^{a}=\psi_{p,h}^{a}\psi_{p^{\prime},h}^{a}$, to reduce the parameter space while allowing a sufficiently flexible structure to address potential confounding bias in our primary predictor of interest. When both sets of responses follow a normal linear model with an identity link function, (1) becomes